OpenStudy (anonymous):

Nancy throws a party and puts the names of 10 friends in a hat to determine who should win a prize. One of the friends draws a name fron the hat,and the prize is given to the person whose name is drawn. Find the probability that the friend draws his or her own name

OpenStudy (anonymous):

2/10 =1/5 =.20

OpenStudy (anonymous):

i got the answer wrong

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

surely that's a 1/10 = .10